# GestureCoach: Rehearsing for Engaging Talks with LLM-Driven Gesture Recommendations

### v1

@@ 7.2. Quality of Predicted Gestures
While our model outperforms off-the-shelf LLMs in predicting gesture regions, the low precision scores raises an important question: Are the detected false positive regions truly inappropriate for gestures, or are some still valid for gesturing? It is possible that a predicted region was not gestured by the original speaker but could still serve as a meaningful emphasis point if accompanied by a suitable semantic gesture. To explore this, we conducted a user study to subjectively evaluate the quality of the gesture regions predicted by our emphasis proposal model and the suitability of the gestures selected by our gesture identification model.
@@ 7.2.2. Results
Overall, participants found the gesture regions proposed by the emphasis proposal model appropriate for emphasis in both held-out ($M=5.23,SD=1.59$) and unseen samples ($M=5.56,SD=1.43$). Similarly, they rated the semantic gestures retrieved by the gesture identification model as well-suited for emphasis in both held-out ($M=5.44,SD=1.58$) and unseen samples ($M=5.08,SD=1.55$). To assess the impact of false positives (see Figure  5), we define “valid” false positives as samples with an average rating of higher than 5 in both appropriateness and suitability. These can be considered true positives, as users found them to be sensible emphasis points. We found that 53.3% of false positives in held-out samples and 60% in unseen samples met this criterion. Additionally, 40% of false positives in both cases scored between 3.5 and 5. Such samples can be considered to be “neutral” or slightly positive, implying gesturing at those points would not negatively affect the presentation. Notably, no samples scored below 3, indicating that the model does not suggest completely irrelevant or unsuitable gesture regions.
@@ 8. User Evaluation
We conducted a user study with 10 participants to evaluate the whether rehearsing presentations with GestureCoach can help users positively affect their gesturing while giving the talk. We also evaluate the user experience and usability of the system to understand its strengths and limitations.
@@ 8.2. Participants
Our participants included 4 males and 6 females, who were proficient in English with an average age of 25.42 years. Their expertise in giving talks in front of an audience varied; 1 participant identified as beginner (1-2 talks), 2 as intermediate (3-5 talks), 5 as advanced (5-10 talks), and 2 experts (10+ talks). Each participant spent 1 hour in the study and received 10 Euros for their time.
Figure 6. Comparison of gesture usage and user preference after rehearsing with GestureCoach vs. Notes. GestureCoach led to significantly more unique semantic gesture use and was preferred by 8 of 10 participants, who reported increased confidence after practice and a more engaging final talk.
[…]
Figure (a) shows a bar graph comparing the average number of unique semantic gestures performed after rehearsing with GestureCoach vs. Notes. GestureCoach is significantly higher than Notes. Figure (b) shows a stacked bar graph of preferences between the two rehearsal methods. 8/10 users preferred GestureCoach overall and found that it improved their confidence and helped deliver a more engaging talk as compared to rehearsing with just Notes.
Figure 7. (a) Subjective evaluation of GestureCoach features indicate its usefulness and ease of use for rehearsal (b) Results of subjective comparison between GestureCoach and Notes.
[…]
Figure (a) shows a horizontal bar graph of the rating distribution on a 5 point Likert scale for various features of the GestureCoach System. Most users agree strongly that the features helped them during rehearsal. Figure (b) a vertical bar graph comparing the rating distribution for GestureCoach and Notes for different aspects such as encouraging gestures, providing gesture ideas, speech intonation etc. Most users strongly agree that GestureCoach improved these aspects while Notes did not. Please refer to the text for specific details.
@@ 8.3.1. Gesture Quantity
[…]
@@ 8.3.2. Gesture Diversity
[…]
@@ 8.3.4. Subjective Comparison of Rehearsal Methods
Rehearsal with GestureCoach was the preferred method for most speakers (8/10), enhancing their confidence (8/10) and enabling them to deliver a more engaging talk (8/10) compared to rehearsing with Notes. As shown in Figure  7b, speakers reported that using GestureCoach during rehearsal encouraged them to incorporate gestures more naturally $(M=4,SD=1.15)$, provided ideas for gesture use $(M=4.70,SD=0.48)$, and helped them use meaningful gestures $(M=4.20,SD=0.63)$ at appropriate points $(M=4.1,SD=0.56)$. It also added diversity to the gestures $(M=4.30,SD=0.67)$ during their talk. Additionally, some participants observed improvements in their intonation $(M=3.70,SD=0.82)$ and pacing $(M=3.10,SD=0.87)$ after rehearsing with GestureCoach. However, they noted that rehearsing with Notes offered more flexibility $(M=4,SD=1.33)$ for integrating their own gestures than GestureCoach $(M=3.70,SD=1.06)$.
@@ 9.1.1. Rehearsing with GestureCoach Helps Deliver More Engaging Talks
Feedback from speakers indicates that the gesture recommendations provided by GestureCoach during rehearsal played a key role in encouraging them to do more gestures during their talk. As P1 described, focusing on gestures when rehearsing with just notes was challenging: “[Without GestureCoach] I know I need to move [my hands], but it’s difficult to think about what would be a good gesture, so I end up just doing random gestures”. In contrast, P1 explained that GestureCoach ”helped me reflect on where to emphasize and what kind of gestures I could use… its like someone giving feedback”. The recommendations provided practical examples and demonstrated how to use gestures more intentionally.
While the primary objective of GestureCoach was to enhance users’ gesturing during presentations, some speakers also reported a synergistic effect on their speech delivery. For instance, P7 felt that rehearsing with GestureCoach improved their intonation: ”When its highlighted I also stress that point in my speech such that it matches the gesture”. This positive impact on intonation occurs because effectively performing a gesture with the intended emphasis requires speakers to slow down and emphasize the associated words, allowing the gesture’s intended meaning to be conveyed more effectively. Thus, by encouraging intentional emphasis in both gestures and speech, GestureCoach helped speakers deliver more expressive and engaging talks.
@@ Appendix B Model details
B.1. Prompt for Emphasis Proposal module
The following
@@ B.1. Prompt for Emphasis Proposal module
B.2. Prompt for Gesture Identification module

### v2

@@ 6.2. Quality of Predicted Gestures
While our model outperforms off-the-shelf LLMs in predicting gesture regions, the low precision scores raise an important question: Are the detected false positive regions truly inappropriate for gestures, or are some still valid for gesturing? It is possible that a predicted region was not gestured by the original speaker but could still serve as a meaningful emphasis point if accompanied by a suitable semantic gesture. To explore this, we conducted a user study to subjectively evaluate the quality of the gesture regions predicted by our emphasis proposal model and the suitability of the gestures selected by our gesture identification model.
@@ 6.2.2. Results
Overall, participants found the gesture regions proposed by the emphasis proposal model appropriate for emphasis in both held-out ($M=5.23,SD=1.59$) and unseen samples ($M=5.56,SD=1.43$). Similarly, they rated the semantic gestures retrieved by the gesture identification model as well-suited for emphasis in both held-out ($M=5.44,SD=1.58$) and unseen samples ($M=5.08,SD=1.55$). To assess the impact of false positives (see Figure  5), we define “valid” false positives as samples with an average rating of higher than 5 in both appropriateness and suitability. These can be considered true positives, as users found them to be sensible emphasis points. We found that 53.3% of false positives in held-out samples and 60% in unseen samples met this criterion. Additionally, 40% of false positives in both cases scored between 3.5 and 5. Such samples can be considered to be “neutral” or slightly positive, implying gesturing at those points would not negatively affect the presentation. Notably, no samples scored below 3, indicating that the model does not suggest completely irrelevant or unsuitable gesture regions. These results highlight the importance of subjective evaluation in generative tasks like gesture prediction, where traditional metrics such as F1 score or segmentation accuracy may not fully capture the quality or effectiveness of the model predictions.
@@ 7. User Evaluation
We conducted a user study with 10 participants to evaluate whether rehearsing presentations with GestureCoach can help users positively affect their gesturing while giving the talk. We also evaluate the user experience and usability of the system to understand its strengths and limitations.
@@ 7.2. Participants
Our participants included 4 males and 6 females, who were proficient in English with an average age of 25.42 years. Their expertise in giving talks in front of an audience varied; 1 participant identified as beginner (1-2 talks), 2 as intermediate (3-5 talks), 5 as advanced (5-10 talks), and 2 as experts (10+ talks). Each participant spent 1 hour in the study and received 10 Euros for their time.
Figure 6. Comparison of gesture usage and user preference after rehearsing with GestureCoach vs. Notes. GestureCoach led to significantly more unique semantic gesture use and was preferred by 8 of 10 participants, who reported increased confidence after practice and a more engaging final talk. :
Figure 7. (a) Subjective evaluation of GestureCoach features indicate its usefulness and ease of use for rehearsal, (b) Results of subjective comparison between GestureCoach and Notes.:
@@ 7.3.1. Gesture Quantity
We quantified semantic gestures performed by speakers during their talks. An expert in the gesture domain (co-author) blind to the conditions annotated the recorded talks using the same criteria used for creating the gesture database (Section  4.1). Overall, speakers performed slightly higher number of gestures in the GestureCoach condition $(M=30.30,SD=8.87)$ than in the Notes $(M=25.1,SD=10.16)$ condition, but the difference was not significant. A closer look at speakers’ individual data showed that 8 of 10 speakers performed significantly more gestures $(t(7)=2.99,p=.02)$ after rehearsing with GestureCoach $(M=31.75,SD=8.97)$ than with Notes $(M=21.65,SD=7.52)$. The remaining two users showed an opposite trend, performing more gestures after using Notes $(M=39,SD=7.07)$ than with GestureCoach $(M=24.5,SD=7.77)$. We discuss potential reasons for this in Section  8.2.2 based on users’ qualitative feedback.
@@ 7.3.2. Gesture Diversity
We measure gesture diversity as the number of unique semantic gestures performed by speakers during their talk. As seen in Figure  6a, speakers performed considerably higher number of unique gestures in the GestureCoach condition $(M=18.90,SD=7.21)$ than in the Notes $(M=12.70,SD=5.10)$ condition. A paired samples t-test showed that the difference is significant $(t(9)=3.27,p=.009)$ with a large effect size, Cohen’s $d=1.03$, suggesting that rehearsing with GestureCoach helped speakers add diversity to the gestures they used during their talk.
@@ 7.3.4. Subjective Comparison of Rehearsal Methods
Rehearsal with GestureCoach was the preferred method for most speakers (8/10), enhancing their confidence (8/10) and enabling them to deliver a more engaging talk (8/10) compared to rehearsing with Notes. As shown in Figure  7b, speakers reported that using GestureCoach during rehearsal encouraged them to incorporate gestures more naturally $(M=4,SD=1.15)$ $(t(9)=3.87,p<.01)$, provided ideas for gesture use $(M=4.70,SD=0.48)$ $(W=55.0,z=2.80,p<.01)$, and helped them use meaningful gestures $(M=4.20,SD=0.63)$ $(W=55.0,z=2.80,p<.01)$ at appropriate points $(M=4.1,SD=0.56)$ $(t(9)=6.00,p<.001)$. It also added diversity to the gestures $(M=4.30,SD=0.67)$ $(t(9)=7.79,p<.001)$ during their talk. Additionally, some participants observed improvements in their intonation $(M=3.70,SD=0.82)$ $(t(9)=1.96,p=0.081)$ and pacing $(M=3.10,SD=0.87)$ $(W=6.0,z=1.60,p=0.174)$ after rehearsing with GestureCoach. However, they noted that rehearsing with Notes offered more flexibility $(M=4,SD=1.33)$ for integrating their own gestures than GestureCoach $(M=3.70,SD=1.06)$, although the difference was not significant.
@@ 8.1.1. Rehearsing with GestureCoach Helps Deliver More Engaging Talks
Feedback from speakers indicates that the gesture recommendations provided by GestureCoach during rehearsal played a key role in encouraging them to do more gestures during their talk. As P1 described, focusing on gestures when rehearsing with just notes was challenging: “[Without GestureCoach] I know I need to move [my hands], but it’s difficult to think about what would be a good gesture, so I end up just doing random gestures”. In contrast, P1 explained that GestureCoach ”helped me reflect on where to emphasize and what kind of gestures I could use… it’s like someone giving feedback”. The recommendations provided practical examples and demonstrated how to use gestures more intentionally.
While the primary objective of GestureCoach was to enhance users’ gesturing during presentations, some speakers also reported a synergistic effect on their speech delivery. For instance, P7 felt that rehearsing with GestureCoach improved their intonation: ”When it’s highlighted I also stress that point in my speech such that it matches the gesture”. This positive impact on intonation occurs because effectively performing a gesture with the intended emphasis requires speakers to slow down and emphasize the associated words, allowing the gesture’s intended meaning to be conveyed more effectively. Thus, by encouraging intentional emphasis in both gestures and speech, GestureCoach helped speakers deliver more expressive and engaging talks.
@@ 8.2.2. Hybrid Rehearsal to Integrate with Natural Gestures
8.2.3. Modeling Considerations for Gesture Recommendation
We approached the gesture recommendation problem as two distinct tasks: deciding when to gesture and determining what gesture to use. The when component was driven by observable and well-established patterns such as gestures frequently occurring at topic changes, during list enumerations, or at points of emphasis. These patterns make it easier to model it as a finetuning task. In contrast, the what component was more complex due to high variability. Semantic gestures follow a long-tail distribution, where a single gesture can suit many different spoken contexts. For such problems, a RAG-based approach may be a better choice as it enables the model to retrieve grounded examples during inference rather than relying entirely on mappings learned through direct fine-tuning.
@@ 8.3.2. Guidance for subtleties in gestures
8.3.3. Design Space of Gesture Visualizations
In our system, we visualized gesture cues using an animated avatar that demonstrated the gesture motion. However, it is also possible to consider other lightweight alternatives, such as emojis or icons, that are embedded within the presentation script. Such cues maybe easier to reference while speaking and could be particularly helpful in real-time contexts. Future work should explore the design space of gesture visualizations and empirically compare them to understand the trade-offs between gesture comprehension and cognitive load.
8.3.4. Adapting to cultural contexts
@@ 9. Conclusion
Acknowledgements.
This work has been partly supported by the Deutsche Forschungsgemeinschaft DFG under Grant No. 521601028 within the Priority Program SPP 2199 Scalable Interaction Paradigms for Pervasive Computing Environments, and the Deutsche Forschungsgemeinschaft, Funder Id: http://dx.doi.org/10.13039/501100001659, SFB 1102: “Information Density and Linguistic Encoding”, project number 232722074. We also thank the reviewers for their insightful comments that helped to improve the paper.
@@ Appendix B Model details
B.1. System Prompt for Emphasis Proposal module
@@
Example Prompt: ¡s¿[INST] ¡¡SYS¿¿ You are an expert in public speaking. The provided Text will be used for a TED talk. Identify the Emphasis Areas in the Text that should emphasized when giving the TED talk.The extracted text in the Emphasis Areas should exactly match phrases in the given Text. ¡¡/SYS¿¿ Text: Cortisol is toxic, and it causes cloudy thinking. So part of the practice of the pre-mortem is to recognize that under stress you’re not going to be at your best, and you should put systems in place. And there’s perhaps no more stressful a situation than when you’re confronted with a medical decision to make. And at some point, all of us are going to be in that position, where we have to make a very important decision about the future of our medical care or that of a loved one, to help them with a decision. And so I want to talk about that. [/INST]
B.2. System Prompt for Gesture Identification module
@@
Example Prompt: You are an expert in public speaking. You are given a TEXT and a QUERY phrase within the text that needs to be emphasized. Additionally, you are provided with CANDIDATE emphasis areas” Your task is to identify the CANDIDATE emphasis phrase that is most similar to the QUERY phrase in the TEXT in terms of emphasis context. Always choose only from the provided list of CANDIDATES. TEXT: First of all, thank you for your attention. There’s nothing quite like being in a room full of people like this, where all of you are giving your attention to me. It’s a powerful feeling, to get attention. I’m an actor, so I’m a bit of an expert on, well, nothing, really. But I do know what it feels like to get attention – I’ve been lucky in my life to get a lot more than my fair share of attention.; QUERY: nothing; CANDIDATES: 1) there’s nothing 2) no matter, 3) the only thing. Give only the candidate as output.’